OpenStudy (abhisar):

Three unequal resistors in parallel are equivalent to a resistance 1 ohm. If two of them are in the ratio 1:2 and if no resistance value is fractional, the largest of three resistance in ohms is

OpenStudy (abhisar):

@ganeshie8

OpenStudy (abhisar):

@Vincent-Lyon.Fr

OpenStudy (anonymous):

answer is 6?

OpenStudy (abhisar):

yes..How ?

OpenStudy (anonymous):

1/R=1/R1+1/2R3+1/R3 ------>1-(3/2-R3)=1/R1--->R1=2R3/2R3-3 If R3 = 3 then 1 R = Ω2 1 ∴R = Ω2 2 R = Ω6 R3 = Ω3 ∴Largest resistance = 6Ω HOPE IT HELPS A BIT.

OpenStudy (anonymous):

R1=2, R2=6, R3=3, HOPE YOU CAN GET IT

OpenStudy (anonymous):

ALL YOU NEED IS FORMULA THAT'S ALL

OpenStudy (abhisar):

\(\sf \frac{1}{R}=\frac{1}{R1}+\frac{1}{2R3}+\frac{1}{R3}\)

OpenStudy (abhisar):

Yes..actually i am a Bio base student so not that good in maths..can u explain me how u deduced the above equation ?

OpenStudy (anonymous):

IT'S EASY, 1/R=1/R1+(3/2R3) FROM THERE KUST YOU NEED TO FIND R1

OpenStudy (anonymous):

R VALUE ALREADY GIVEN AND RATIO IS MENTIONED 1:2

OpenStudy (abhisar):

ok so ratio is given 1:2..and u hav supposed R1:R3 = 1:2 ?

OpenStudy (anonymous):

NO, ITS R2/R3=1:2

OpenStudy (abhisar):

ok...i got that but can u explain this If R3 = 3 then 1 R = Ω2

OpenStudy (abhisar):

y u supposed R3 as 3 ?

OpenStudy (abhisar):

Let the three resistors be y, x and 2x So \(\sf \huge 1=\frac{1}{x}+\frac{1}{2x}+\frac{1}{y}\\ \Rightarrow \huge y=\frac{2x}{2x-3}\\ \text{Now x=0 and x=1 is not possible because it will give the value of y as 0 and -2}\\ \text{Any value >3 wil make the denominator>Numerator and thus y will not be integer}\\ \text{So only possible values left are 2 and 3. Since we want the value to be greatest}\\ \text{we take x=3, which give the values 6,3 and 2}\)

OpenStudy (anonymous):

true...!!!

OpenStudy (abhisar):

Thank you !

OpenStudy (anonymous):

np)